How many 4-digit numbers can you build with a 3, a 4, a 6, and an 8 that are over 6000?
Answer by stanbon(75887)

How many 4-digit numbers can you build with a 3, a 4, a 6, and an 8 that are over 6000?
---
The thousands digit would have to be 6 or 8, i.e. 2 ways
---
Each of the remaining 3 digits can be selected
in 4 ways, if repetition is allowed.
---
That would give you 2*4^3 = 2*64 = 128 numbers
=================
If repetition is not allowed you get 2*3*2*1 = 12 numbers.
=================================================================
Cheers,
Stan H.